# The Belle II Collaboration

### v2

@@ Abstract
Figure 3: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % confidence level from PS191 [55], E949 [56], NA62 [57, 58], KOTO [59, 16], KTeV [60], and BABAR [61, 54], and at 95 % confidence level from MicroBooNE [62, 63, 16], L3 [64, 16], CHARM [65, 54], LHCb [66, 67, 54], Belle II  [68], and CMS [69] for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, $\Delta m=0.4\,m(\chi_{1})$, $\epsilon=1.5\times 10^{-3}$, and $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints except for the one from this work do not require the presence of a dark photon or iDM.
@@
Figure 4: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % confidence level from CHARM [23, 18], NuCal [21, 22, 18], and BABAR [20, 12]) for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, $\Delta m=0.4\,m(\chi_{1})$, $\sin\theta=2.6\times 10^{-4}$, and $m(h^{\prime})=0.4\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints except for the one from this work do not require the presence of a dark Higgs boson or iDM.
@@ Appendix E Upper Limits
Figure S17: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [55], E949 [56], NA62 [57, 58], KOTO [59, 16], KTeV [60], and BABAR [61, 54], and at 95 % credibility level from MicroBooNE [62, 63, 16], L3 [64, 16], CHARM [65, 54], LHCb [66, 67, 54], Belle II  [68], and CMS [17] for $\alpha_{D}=0.1$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=3\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=0.01\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=0.1\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S18: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [55], E949 [56], NA62 [57, 58], KOTO [59, 16], KTeV [60], and BABAR [61, 54], and at 95 % credibility level from MicroBooNE [62, 63, 16], L3 [64, 16], CHARM [65, 54], LHCb [66, 67, 54], Belle II  [68], and CMS [17] for $\alpha_{D}=0.1$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=3\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=1.0\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=100.0\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. For the model parameter configurations in the center right and lower right plot Belle II is not sensitive. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S19: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [55], E949 [56], NA62 [57, 58], KOTO [59, 16], KTeV [60], and BABAR [61, 54], and at 95 % credibility level from MicroBooNE [62, 63, 16], L3 [64, 16], CHARM [65, 54], LHCb [66, 67, 54], Belle II  [68], and CMS [17]) for $\alpha_{D}=0.1$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=4\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=0.1\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=1.0\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S20: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [55], E949 [56], NA62 [57, 58], KOTO [59, 16], KTeV [60], and BABAR [61, 54], and at 95 % credibility level from MicroBooNE [62, 63, 16], L3 [64, 16], CHARM [65, 54], LHCb [66, 67, 54], Belle II  [68], and CMS [17] for $\alpha_{D}=0.5$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=3\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=0.01\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=1.0\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S22: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % credibility level from CHARM [23, 18], NuCal [21, 22, 18], and BABAR [20, 12] for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, and $m(h^{\prime})=0.6\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top) and $\Delta m=0.4\,m(\chi_{1})$ (bottom). The mixing angle of the dark Higgs is set to $\sin\theta=7.0\times 10^{-4}$ ($c\tau(h^{\prime})=1.0\,{\mathrm{\,cm}}$) (left) and $\sin\theta=1.5\times 10^{-4}$ ($c\tau(h^{\prime})=21.54\,{\mathrm{\,cm}}$) (right). Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints but the one from this work do not depend on the presence of a dark Higgs boson or iDM.
Figure S23: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % credibility level from CHARM [23, 18], NuCal [21, 22, 18], and BABAR [20, 12] for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, $m(h^{\prime})=1.2\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $\sin\theta=8.8\times 10^{-5}$ ($c\tau(h^{\prime})=21.54\,{\mathrm{\,cm}}$). The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (left) and $\Delta m=0.4\,m(\chi_{1})$ (right). Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints but the one from this work do not depend on the presence of a dark Higgs boson or iDM.
Figure S24: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % credibility level from CHARM [23, 18], NuCal [21, 22, 18], and BABAR [20, 12] for $\alpha_{D}=0.1$, $\Delta m=0.4\,m(\chi_{1})$, $m(h^{\prime})=0.4\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $\sin\theta=2.6\times 10^{-4}$ ($c\tau(h^{\prime})=21.54\,{\mathrm{\,cm}}$). The dark photon mass is set to $m(A^{\prime})=3\,m(\chi_{1})$ (left) and $m(A^{\prime})=4\,m(\chi_{1})$ (right). Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints but the one from this work do not depend on the presence of a dark Higgs boson or iDM. For $m(A^{\prime})=4\,m(\chi_{1})$ no recasts of the BaBar, CHARM, and NuCal exclusion regions exist in the literature.

### v3

@@ Abstract
Figure 3: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % confidence level from PS191 [Gorbunov:2021ccu], E949 [BNL-E949:2009dza], NA62 [NA62:2020pwi, NA62:2021zjw], KOTO [KOTO:2020prk, Ferber:2023iso], KTeV [KTEV:2000ngj], and BABAR [BaBar:2015jvu, Winkler:2018qyg], and at 95 % confidence level from MicroBooNE [MicroBooNE:2021usw, MicroBooNE:2022ctm, Ferber:2023iso], L3 [L3:1996ome, Ferber:2023iso], CHARM [CHARM:1985anb, Winkler:2018qyg], LHCb [LHCb:2015nkv, LHCb:2016awg, Winkler:2018qyg], Belle II[Belle-II:2023ueh], and CMS [CMS:2021sch] for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, $\Delta m=0.4\,m(\chi_{1})$, $\epsilon=1.5\times 10^{-3}$, and $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints except for the one from this work do not require the presence of a dark photon or iDM.
@@
Figure 4: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % confidence level from CHARM [Gninenko:2012eq, Tsai:2019buq], NuCal [Blumlein:2011mv, Blumlein:2013cua, Tsai:2019buq], and BABAR [BaBar:2017tiz, Duerr:2019dmv]) for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, $\Delta m=0.4\,m(\chi_{1})$, $\sin\theta=2.6\times 10^{-4}$, and $m(h^{\prime})=0.4\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints except for the one from this work do not require the presence of a dark Higgs boson or iDM.
@@ Appendix E Upper Limits
Figure S17: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [Gorbunov:2021ccu], E949 [BNL-E949:2009dza], NA62 [NA62:2020pwi, NA62:2021zjw], KOTO [KOTO:2020prk, Ferber:2023iso], KTeV [KTEV:2000ngj], and BABAR [BaBar:2015jvu, Winkler:2018qyg], and at 95 % credibility level from MicroBooNE [MicroBooNE:2021usw, MicroBooNE:2022ctm, Ferber:2023iso], L3 [L3:1996ome, Ferber:2023iso], CHARM [CHARM:1985anb, Winkler:2018qyg], LHCb [LHCb:2015nkv, LHCb:2016awg, Winkler:2018qyg], Belle II[Belle-II:2023ueh], and CMS [CMS:2023bay] for $\alpha_{D}=0.1$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=3\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=0.01\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=0.1\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S18: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [Gorbunov:2021ccu], E949 [BNL-E949:2009dza], NA62 [NA62:2020pwi, NA62:2021zjw], KOTO [KOTO:2020prk, Ferber:2023iso], KTeV [KTEV:2000ngj], and BABAR [BaBar:2015jvu, Winkler:2018qyg], and at 95 % credibility level from MicroBooNE [MicroBooNE:2021usw, MicroBooNE:2022ctm, Ferber:2023iso], L3 [L3:1996ome, Ferber:2023iso], CHARM [CHARM:1985anb, Winkler:2018qyg], LHCb [LHCb:2015nkv, LHCb:2016awg, Winkler:2018qyg], Belle II[Belle-II:2023ueh], and CMS [CMS:2023bay] for $\alpha_{D}=0.1$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=3\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=1.0\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=100.0\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. For the model parameter configurations in the center right and lower right plot Belle II is not sensitive. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S19: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [Gorbunov:2021ccu], E949 [BNL-E949:2009dza], NA62 [NA62:2020pwi, NA62:2021zjw], KOTO [KOTO:2020prk, Ferber:2023iso], KTeV [KTEV:2000ngj], and BABAR [BaBar:2015jvu, Winkler:2018qyg], and at 95 % credibility level from MicroBooNE [MicroBooNE:2021usw, MicroBooNE:2022ctm, Ferber:2023iso], L3 [L3:1996ome, Ferber:2023iso], CHARM [CHARM:1985anb, Winkler:2018qyg], LHCb [LHCb:2015nkv, LHCb:2016awg, Winkler:2018qyg], Belle II[Belle-II:2023ueh], and CMS [CMS:2023bay]) for $\alpha_{D}=0.1$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=4\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=0.1\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=1.0\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S20: Exclusion regions at 95 % credibility level in the plane of the sine of the mixing angle $\theta$ and dark Higgs mass $m(h^{\prime})$ from this work (teal) together with existing constraints at 90 % credibility level from PS191 [Gorbunov:2021ccu], E949 [BNL-E949:2009dza], NA62 [NA62:2020pwi, NA62:2021zjw], KOTO [KOTO:2020prk, Ferber:2023iso], KTeV [KTEV:2000ngj], and BABAR [BaBar:2015jvu, Winkler:2018qyg], and at 95 % credibility level from MicroBooNE [MicroBooNE:2021usw, MicroBooNE:2022ctm, Ferber:2023iso], L3 [L3:1996ome, Ferber:2023iso], CHARM [CHARM:1985anb, Winkler:2018qyg], LHCb [LHCb:2015nkv, LHCb:2016awg, Winkler:2018qyg], Belle II[Belle-II:2023ueh], and CMS [CMS:2023bay] for $\alpha_{D}=0.5$, $m(\chi_{1})=2.5\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $m(A^{\prime})=3\,m(\chi_{1})$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top), $\Delta m=0.4\,m(\chi_{1})$ (center), and $\Delta m=1.0\,m(\chi_{1})$ (bottom). Plots on the left assume a $\chi_{2}$ lifetime of $c\tau(\chi_{2})=0.01\,{\mathrm{\,cm}}$ and on the right $c\tau(\chi_{2})=1.0\,{\mathrm{\,cm}}$. This results in different mixing parameters $\epsilon$, which are reported in each plot. All constraints but the one from this work do not depend on the presence of a dark photon or iDM.
@@
Figure S22: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % credibility level from CHARM [Gninenko:2012eq, Tsai:2019buq], NuCal [Blumlein:2011mv, Blumlein:2013cua, Tsai:2019buq], and BABAR [BaBar:2017tiz, Duerr:2019dmv] for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, and $m(h^{\prime})=0.6\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$. The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (top) and $\Delta m=0.4\,m(\chi_{1})$ (bottom). The mixing angle of the dark Higgs is set to $\sin\theta=7.0\times 10^{-4}$ ($c\tau(h^{\prime})=1.0\,{\mathrm{\,cm}}$) (left) and $\sin\theta=1.5\times 10^{-4}$ ($c\tau(h^{\prime})=21.54\,{\mathrm{\,cm}}$) (right). Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints but the one from this work do not depend on the presence of a dark Higgs boson or iDM.
Figure S23: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % credibility level from CHARM [Gninenko:2012eq, Tsai:2019buq], NuCal [Blumlein:2011mv, Blumlein:2013cua, Tsai:2019buq], and BABAR [BaBar:2017tiz, Duerr:2019dmv] for $\alpha_{D}=0.1$, $m(A^{\prime})=3\,m(\chi_{1})$, $m(h^{\prime})=1.2\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $\sin\theta=8.8\times 10^{-5}$ ($c\tau(h^{\prime})=21.54\,{\mathrm{\,cm}}$). The mass splitting is set to $\Delta m=0.2\,m(\chi_{1})$ (left) and $\Delta m=0.4\,m(\chi_{1})$ (right). Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints but the one from this work do not depend on the presence of a dark Higgs boson or iDM.
Figure S24: Exclusion regions at 95 % credibility level in the plane of the dimensionless variable $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ and DM mass $m(\chi_{1})$ from this work (teal) together with existing constraints at 90 % credibility level from CHARM [Gninenko:2012eq, Tsai:2019buq], NuCal [Blumlein:2011mv, Blumlein:2013cua, Tsai:2019buq], and BABAR [BaBar:2017tiz, Duerr:2019dmv] for $\alpha_{D}=0.1$, $\Delta m=0.4\,m(\chi_{1})$, $m(h^{\prime})=0.4\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$, and $\sin\theta=2.6\times 10^{-4}$ ($c\tau(h^{\prime})=21.54\,{\mathrm{\,cm}}$). The dark photon mass is set to $m(A^{\prime})=3\,m(\chi_{1})$ (left) and $m(A^{\prime})=4\,m(\chi_{1})$ (right). Constraints colored in gray with dashed outline are reinterpretations not performed by the experimental collaborations. All constraints but the one from this work do not depend on the presence of a dark Higgs boson or iDM. For $m(A^{\prime})=4\,m(\chi_{1})$ no recasts of the BaBar, CHARM, and NuCal exclusion regions exist in the literature.